### v1

# WebQuest: A Benchmark for Multimodal QA on Web Page Sequences

Abstract
The rise of multimodal LLMs and web agents calls for the creation of challenging benchmarks to evaluate neural architectures. Unlike existing benchmarks that focus on multi-step web navigation, we present WebQuest, a multi-page question-answering dataset that requires simultaneous retrieval and reasoning across web interaction sequences grounded in real-world usage. WebQuest includes three question categories: single-screen reasoning, multi-screen reasoning, and questions based on navigation traces. We evaluate some of the leading multimodal models like GPT-4V, Gemini Flash, and Claude 3 on our dataset, revealing a significant gap between single-screen and multi-screen reasoning. Finally, we investigate inference time techniques like Chain-of-Thought prompting to improve model capabilities on multi-screen reasoning.
WebQuest: A Benchmark for Multimodal QA on Web Page Sequences
Maria Wangthanks: First authors. Correspondence: jdchen@google.com, Srinivas Sunkara11footnotemark: 1, Gilles Baechlerthanks: Equal contribution, Jason Lin22footnotemark: 2, Yun Zhu, Fedir Zubach, Lei Shu, Jindong Chen Google DeepMind
1 Introduction
11todo: 1jdchen: last pass22todo: 2jasonjlin: done
Figure 1: An example of multi-page question and answering in WebQuest. In this examples, there are 3 screenshots and 2 question-answer pairs. It requires information extraction from each screenshots and reasoning in order to answer both questions.
As the de facto I/O for Human-Computer Interaction, Graphical User-Interfaces (UIs) allow users to acquire information and complete tasks similar to APIs, as the native interface between computer software. A predominant type of data we consume are multimodal web pages rich in semantic knowledge and structure. The advent of smart devices has promised humans the ease of use, motivating research for multimodal, assistive, predictive, automatic and personalized technology for smart UIs. However, designers of such systems do not always consider that the Internet is modeled after our cognitive patterns, such as a limited attention span and working memory. Bounded by single-page interfaces and the ability to traverse back and forward in a sequence, machine learning techniques have been proposed for UI automation but success on multi-step, multi-screen settings are limited, making them unable to fulfill sophisticated tasks for users. Additionally, there are very few datasets focused on information extraction across navigation traces versus task completion.
Recently, Multimodal Large Language Models (MLLMs) like Gemini Team et al. (2023), GPT-4V OpenAI et al. (2024) have been trained on text and image data crawled from the web to demonstrate powerful zero shot capabilities. However, the data collection process of visual-language datasets Schuhmann et al. (2021, 2022)33todo: 3Which citation(s) we have in mind here? reveals a discrepancy with how a human user would interact with the web: navigating multiple pages searching for information using the various UI elements present on the UI. By simply extracting text and images from web pages, these datasets not only remove the functional congruence shared between desktop and mobile form factors, but also fail to capture the actions taken and the implicit retrieval requirement that is common to a information-seeking or goal-oriented workflow, i.e. flight booking or shopping. As a result, these models are effective in general reasoning, but perform poorly when used as end-to-end autonomous agents.
A key challenge in developing UI agents is that they not only have to interpret spatial relations between structures, they also have to ground open ended instructions within an expanding set of semantically correlated pages across diverse applications. Traditional NLP benchmarks like question answering (QA) and information retrieval have been adapted for visual grounding, but they contain either single-page or multi-page questions that contain the answer on one of the pages. We postulate the lack of a temporal dimension might stem from human’s limited working memory and the limitations of existing models to effectively predict such complex interactions. When provided a sequence of pages, local similarity between adjacent pages requires a model to discern relevant information among confounding concepts. It can therefore be limiting to develop and evaluate LLM agents using only the standard QA formulation. To this end, we modify the QA formulation to include multiple pages, which more accurately reflects a user’s ability to compare and reason across pages, i.e. “If I restrict the take-off time to be between 8:00 AM and 8:00 PM, how much more I need to pay for the best choice in business class?” to arrive at an informed decision or conclusion. We do acknowledge that some sophisticated websites already provide filters and comparison functionalities, but they are limited to popular options and do not always expose all the functionality available to the user.
Motivated by this question, we introduce WebQuest, a multi-page screen question-answering dataset, to emulate both navigation and retrieval reasoning settings formulated as a VQA evaluation of multimodal web agents. The dataset and related code will be released on github at: <to-be-updated>. Figure 1 illustrates an example of multi-screen question answering. We seek to close the gap between general QA-based content understanding and agent model research by providing a novel mode of QA setting. Our dataset targets question-answering across web interaction sequences grounded in real world user behavior. WebQuest includes three question categories: single-screen QA, multi-screen QA, and questions based on navigation traces. The first two involve multimodal content from one or more screens, while navigation trace questions are based on sequences of screens navigated by humans to perform routine tasks. This necessitates the endowment of retrieved knowledge from relevant parts of the sequence to aid with complex screen understanding.
Finally, we evaluate state-of-the-art models GPT-4V OpenAI et al. (2024), Gemini Flash Team (2024), and Claude 3 Anthropic (2024) via prompt engineering and Chain-of-thought prompting, revealing a significant gap between single-page and multi-page reasoning. The main contributions of our work are multifold and we believe it will advance the field of UI understanding and automation.
  • WedQuest is the first dataset containing QA over multiple screens and navigation traces requiring information from more than one screen to answer the question.
  • We release 3 subsets of data containing single-screen, multi-screen and trace images allowing more detailed benchmarking of the capabilities of different models
  • For the trace QA split, we also provide step-by-step supervision for the data which can help improve models on navigation with the aim of answering questions.
2 Related Work
44todo: 4jasonjlin: done
WebQuest is a dataset for Visual Question Answering across multiple images on web screens. In addition, it also includes navigation traces composed of the sequence of steps taken to arrive at the relevant screens to answer the questions. We present prior research along the different aspects covered by our dataset below.
Visual Question Answering
The problem of answering questions based on the contents of an image has been studied extensively by the research community from natural image VQA Antol et al. (2015), to TextVQA Singh et al. (2019) of text in natural images and DocVQA Mathew et al. (2021) containing documents with text, tabular structures and figures. Focusing more on documents and UIs, datasets like ChartQA Masry et al. (2022) and InfographicVQA Mathew et al. (2022) emphasize more complex numerical and compositional reasoning. For web and UI understanding, QA datasets like WebSRC Chen et al. (2021) focused on desktop screens and ScreenQA Hsiao et al. (2022) focused on mobile app screens contain questions which require extracting relevant information from a single screen. These two datasets are representative of related work. However, our dataset contains questions that necessitate more complex reasoning and also span multiple screens. Similar to our dataset in having multiple images as input, Multi-page DocVQA Tito et al. (2023) requires visual extraction from up to 20 pages to answer questions. However, Multi-page DocVQA requires identifying a single page that contains the answer whereas in our dataset, the answer spans multiple screens. To the best of our understanding, there are no QA benchmarks spanning multiple websites or images in a task-oriented sequence, where a subset of pages are necessary for joint reasoning.
Visually-situated Grounding
Computer control and workflow automation involve decision-making grounded in visual inputs. Agents have to plan across diverse APIs, interfaces to execute the correct action at the OS and application level. Datasets either captures grounding in a multi-task setting via natural language expressions Liu et al. (2024) or end-to-end Rawles et al. (2024). Previous modeling focused on text-based observations, while pixel-inputs are increasingly explored in recent work like Shaw et al. (2023), Yang et al. (2023), and Zheng et al. (2024). Despite the increase in datasets related to single-page action grounding and multi-step navigation as discussed in the section below, our dataset is the first to focus on question answering in a single and multi page setting.
Web Navigation and Agent Benchmarks
Recently datasets that focus on performing various tasks in web and mobile scenarios have garnered widespread attention. Earlier efforts introduced simulated web and mobile environments, such as MiniWob++ Liu et al. (2018), MoTIF Burns et al. (2022), Mind2Web Deng et al. (2023) and WebShop Yao et al. (2023). To facilitate autonomous web agents, WebArena Zhou et al. (2024), VisualWebArena Koh et al. (2024) and WebLinx Lù et al. (2024) pair interactive instructions with DOM trees, HTML and pixel-based environments. A majority of these datasets do not have step-by-step supervision for each action taken as part of completing the task. Recently released MMInA Zhang et al. (2024) is the first to collect multi-hop, sequential navigation across websites. While they focus on long-range reasoning with only one website per task category (e.g. shopping, travel), WebQuest has multiple websites for each task which is closer to real world workflows.
3 WebQuest Benchmark
Question: How many more days will it be cloudy than rainy?
Answer: 1
Figure 2: An example of Single Screen QA, where the task is to count how many cloudy days before a rainy one, and the weather conditions are depicted by pictograms.
Question: How much do all the gym accessories cost in total?
Answer: \$127.77
Figure 3: An example of Multi Screen QA, about prices of selected gym accessories, each from a different online shop. Notice the layouts and font sizes vary among them.
WebQuest is a multimodal benchmark for question answering in a web navigation setting. It is targeted to replicate the scenarios where users navigate to websites to retrieve information across these different websites. To better study the capabilities and failure cases of various modeling approaches, we split these scenarios into three categories:
  1. 1.
    Single screen QA: Questions about the web page the user is currently on. (see Figure 2 for an example);
  2. 2.
    Multi screen QA: Questions across multiple web pages simultaneously displayed to the user in their browser, where the web pages are shortlisted such that all of them are required to answer the question. (see Figure 1 and Figure 3 for examples);
  3. 3.
    Trace QA: Questions about a user’s browsing session containing the entire sequence of screens viewed by the user. (see Figure 13 in the Appendix for an example).
To build a benchmark that represents these three scenarios, we render and capture screenshots of single pages, sets of 3 to 5 relevant pages, and browsing sessions with on average 16 screens, respectively. The questions in this benchmark all involve logical reasoning, requiring models to analyze screen images, extract relevant information from each page, and finally combine the information across pages to answer the question. In this section, we go into details of each of the categories of WebQuest benchmark and discuss the data collection process and then we present detailed statistics of our dataset.
3.1 Single screen Question Answering
The single screen QA benchmark consists of 542 web screenshots with a question-answer pair associated with each screen. To collect the screenshots and associated QA pairs, we provide the raters a list of most popular websites. Then we ask the raters to browse these popular websites, consider the kind of questions a user navigating this website might be interested in, and take a screenshot and generate the relevant QA pair for each such page. The QA pair is required to involve reasoning, including arithmetic (counting, addition, subtraction) and comparison questions. Additionally, we ask the raters to ensure that the questions are diverse and reflect real world scenarios. To address safety and privacy concerns, the raters are asked to only navigate to safe web pages without inappropriate content. They are also instructed to not sign in to any of the websites. Lastly, we do a round of screenshot validations to filter out unsafe screens and ones containing PII. In Figure 4, we present the distribution of different website categories.
Figure 4: Distribution of website categories of Single Screen QA examples.
3.2 Multi screen Question Answering
For multi screen question answering benchmark, we provide 171 QA pairs each containing multiple screens. In the multi screen setting, we aim to capture the case where a user has navigated to several relevant pages simultaneously, most commonly in different tabs, then want to get answers to questions about these different web pages, such as calculating prices of products across websites and comparing different product attributes. To ensure the screens in the same example are relevant, we construct a list of popular websites and group them into different categories, such as shopping, flights, etc. We provide these websites lists to human raters and they are required to navigate to 3 to 5 websites listed in the same category. For example, these sets can be menu pages of several restaurants or flight search pages across different booking websites. After navigating to the web pages, the raters generate the associated question-answer pair with the example. Similar to the single screen questions, the questions are focused on logical reasoning, such as comparisons, counting, filtering, arithmetic, etc. In Figure 5, we present the distribution of number of screens per example. 55todo: 5[Non Urgent]To Maria, for each graph you created in this section, please add a link to the colab or source for it, for future reference.
Figure 5: Distribution of screen lengths of Multi Screen QA examples.
3.3 Trace Question Answering
The trace question answering benchmark consists of 292 examples. The screens in this category represents a sequence of web pages the user navigates to while obtaining the answer to a question. In all cases, only a subset of the screens are relevant to answering the question. As the number of screens increases in the navigation trace, to answer questions based on the screens seen, a representation of the relevant information encountered needs to be maintained. The main difference between this category and the Multi screen QA category is that, in Multi Screen QA each QA pair contains only the screens with information relevant to answer the question, whereas in Trace QA each QA pair also contains the different screens encountered in the same browsing session. We consider Multi screen QA as a simplified version of the Trace QA task which enables us to measure the capabilities of different modeling approaches on answering questions across multiple screens.
For the data collection, similar to the multi-screen screen QA data collection, we provide top websites and their associated categories to the human raters. The raters browse through websites and record screen traces. On average, each trace contains 16 screens. We display the distribution of the number of screens in Figure 6, and the distribution in different categories in Figure 7. Please see Table 4 in the Appendix for exact counts of the websites for the different questions. After constructing the screen navigation history, the rater reviews the screens and generates QA pairs. The raters are instructed to create questions involving multiple screens in the screen sequence and requiring logical or arithmetic reasoning.
Figure 6: Distribution of the length of traces for the Trace QA benchmark. There are a few very long traces with more than 50 screens, while a majority of the traces contain fewer than 20 screens.
Figure 7: Distribution of categories of Trace QA examples. We observe that news, transportation and shopping categories contain the most number of traces.
4 Experiments
inline,color=redinline,color=redtodo: inline,color=redCan you check this?
4.1 Metrics
To evaluate the performance of various models on the WebQuest benchmark, we utilize a variant of the Relaxed Accuracy metric used for the ChartQA Masry et al. (2022) dataset. It is noteworthy that over half of the answers in this benchmark are numerical values. For these numerical values answers, we use heuristics to extract the numbers from predictions and ground truths. For floating-point answers, similar to Relaxed Accuracy. we allow a margin of ±0.05 in predictions to be considered correct. In contrast, for integer and string answers, we use SQuaD Rajpurkar et al. (2016) pre-processing and exact match between the prediction and ground truth as the success criterion. Furthermore, we incorporate multiple ground truth variants generated using Gemini Pro Team (2024) to accommodate various acceptable answer formats and consider the final score as the maximum score across all ground truth variants. 66todo: 6sinivasksun, jasonjlin inline,color=redinline,color=redtodo: inline,color=redsri:Did we add any evals for few shot prompting? If not, we should remove the few shot prompting mentions?
4.2 Baselines
We evaluate the performance of a number of state-of-the-art MLLMs, along with the adapted forms of web-oriented models on the WebQuest benchmark. We use pretrained parameters and default settings of API-based models for evaluation. The model input includes the screen image(s) and question. Although image ordering should not be relevant to the questions, we use the same ordering of images for Multi Screen QA and the original navigation sequence orders for Trace QA. We craft the model prompts to only output only numbers for Single and Multi Screen QA and numbers or short phrases for Trace QA. For models we used out of the box, inference time techniques like prompt engineering and Chain-of-thought prompting Wei et al. (2022) were utilized to improve their performance. In our Chain-of-thought Wei et al. (2022) prompt, we include stages to analyze the screenshot and question, extract information from each screen, analyze extracted information, then finally to generate the answer. We evaluated various state of the art models like GPT-4V OpenAI et al. (2024), Gemini Flash Team (2024), Claude 3 Anthropic (2024). For Claude 3 Anthropic (2024) models, the model input is limited to 20 images. Therefore for Trace QA evaluations, we only use the last 20 images in the sequence. Furthermore, due to memory constraints for Claude 3 Anthropic (2024) models, we resize the images preserving aspect ratio such that the shorter side is 1440 pixels for Trace QA, whereas we use the original image resolutions for Single screen QA and Multi screen QA. For GPT-4V OpenAI et al. (2024) and Gemini Flash Team (2024) evals we use the original image resolutions for all questions.
Model SingleScreen QA MultiScreen QA Trace QA
no CoT CoT no CoT CoT no CoT CoT
Claude-3 Sonnet 27.3 38.9 13.9 41.9 23.2 19.6
Claude-3 Opus 25.3 40.4 15.1 23.3 18.5 26.6
Gemini Flash 34.9 48.9 11.1 46.8 24.6 35.9
GPT4-V 36.7 57.4 24.1 46.5 29.0 40.8
Table 1: Performance of various MLLMs on variants of the WebQuest benchmark. All these models are used in a zero shot setting without any fine-tuning. For each of the models, prompt tuning was performed separately. Overall, we can see that all of the MLLMs perform a lot better on the Single Screen QA split and their performance decreases as the number of screens increases. Also, note that the performance of the MLLMs on the Trace QA split is considerably lower than on the Multi Screen QA split.
Figure 8: Performance of different models vs number of screens in each question. Perhaps surprisingly, we do not see clear evidence of model performance reducing as number of screens increases.
Figure 9: Performance of different models for different question types. The types ending in a + sign e.g., arithmetic+ indicate questions which contain multiple types with arithmetic being one of them. Here we notice that the different models do not perform well on questions involving counting, whereas they do well on arithmetic questions. All models presented here were evaluated with CoT prompts.
4.3 Main Results
In this section, we present a comprehensive comparison of different MLLMs on our WebQuest benchmark. To better understand the capabilities and failure cases of different models we present results on the various subsets described in section 3. The best performance for each MLLM on the WebQuest splits can be seen in table 1. We analyze the results in the sections below.
4.3.1 Single Screen QA
As described in section 3, the single screen question answering split involves answering a question based on the content of a single web screen. On this task, we notice that all the MLLMs perform relatively well compared to Multi Screen QA and Trace QA. Chain-of-thought Wei et al. (2022) prompting improves the scores significantly. We also note that the performance difference for large MLLMs is very close on this split when compared to a model like ScreenAI Baechler et al. (2024) specifically trained on UI data.
4.3.2 Multi Screen QA
To recap briefly, the multi screen question answering split involves question answering across various relevant screens from different websites. As we saw in section 3, the images are from related websites and on average, each QA pair contains 3.84 screens with 3, 4 or 5 screens. For this split, we notice that all the MLLMs perform worse than the single screen split. With some MLLMs, we notice considerable performance improvement with Chain-of-thought Wei et al. (2022) prompting and few shot evaluations. However, the performance for all the MLLMs drops to around 45% on this split, highlighting the difficulty some of these models have in reasoning across multiple images. We further analyze the performance gap in section 4.4.
4.3.3 Trace QA
The main difference between this split and the Multi Screen QA 4.3.2 is that this split also contains the screens seen during navigation to the relevant screens for answering the questions. As stated in section 3, the Multi Screen QA split is a simplified version of this split as it already contains only the relevant screens. This split can also be used for evaluating different agent systems since it contains action information for each step enabling step-by-step supervision compared to existing datasets. On average, each question has $15.815.8$ 77todo: 7Update this number from the stats screens, out of which relevant information is present on $3.23.2$ color=redcolor=redtodo: color=redUpdate this number from the stats - how do we compute this? screens.
We notice that the performance of all the models is significantly worse on this split compared to both the Single Screen and Multi Screen QA splits. With Chain-of-thought prompting Wei et al. (2022) and prompt engineering, MLLMs on average achieve a score of 30.6%. Out of the MLLMs we evaluated, GPT-4V OpenAI et al. (2024) achieves the best performance with a score of 40.8% on this dataset. Claude 3 Anthropic (2024) models performs significantly worse than GPT-4V OpenAI et al. (2024) and Gemini Flash Team (2024) models, likely due to the extra image resizing step to avoid out of memory errors.
4.4 Analysis
In this section, we compare the performance of different MLLMs on WebQuest in Table 1. We highlight some of our significant findings below.
Single screen vs Multi screen
We observe that all the models including the largest ones achieve much higher performance on single screen tasks compared to multi screen tasks. However, as we can see in figure 8, for the MultiScreen QA split, there is no clear evidence of model performance decreasing as the number of relevant screens increase. This observation needs further investigation. From figure 9 we observe a significant performance gap in questions involving counting across screens compared to arithmetic or comparison operations.
Impact of prompting techniques on results
We observe that prompting techniques like Chain-of-thought Wei et al. (2022), asking the models to split the question into its constituent parts and describing the contents of each screen results in significant improvements across different kinds of models, even with the largest models across the GPT-4 OpenAI et al. (2024), Gemini Team (2024) and Claude 3  Anthropic (2024) model families. On average, we see improvements of 26.5 percentage points on Multi screen QA split and 6.7 percentage points on the Trace QA split. When we manually observe the CoT responses generated, we do note that the model is able to answer sub-questions on various screens correctly and this ability results in better overall performance.
Varying performance on different question categories
For each of the questions in the dataset, we also provide information on the kind of reasoning needed to answer the question. Inspired by similar annotations in the InfographicVQA Mathew et al. (2022) dataset, we use the categories: arithmetic, comparison, counting and filtering where each question can have more than one category. As shown in figure 9, we note that different models exhibit different performances across the various categories. Across all models, questions related to counting tend to get a lower accuracy than other ones, while questions involving only arithmetic operations get a higher score. However, it is important to note that these aggregate numbers relying on the final answer are conflating errors due to reasoning with errors due to information extraction or other causes. We also note that the performance decreases as the number of entities across the screens increases indicating the model may not able to extract all the information correctly when a large number of entities are present.
5 Conclusion
In this work, we introduce WebQuest, the first multimodal question answering benchmark designed to evaluate reasoning abilities across multiple screens. Our benchmark encompasses three distinct settings: single-screen, multi-screen, and trace-based, enabling a comprehensive assessment of model capabilities. Evaluations of state-of-the-art MLLMs, including Gemini Flash Team (2024), Claude 3 Anthropic (2024), and GPT-4V OpenAI et al. (2024), reveal a significant performance gap between single-screen and multi-screen reasoning tasks. Notably, Chain-of-thought prompting Wei et al. (2022) proves effective for navigating complex, multi-screen scenarios by facilitating information extraction and synthesis across multiple screens. Future research directions include leveraging richer screen information, such as OCR, DOM, and screen annotations from tools like ScreenAI Baechler et al. (2024), to potentially enhance model performance and even replace raw screen images in cases of long screen sequences. Furthermore, extending WebQuest to support personalized, multi-turn dialogue scenarios presents a compelling avenue for future exploration, enabling the development of agents capable of assisting users in dynamic, context-aware conversations.
Appendix A Dataset Examples
We show a few examples of the WebQuest dataset, in Figure 10 for Single Screen QA, in Figures 11 and 12 for Multi Screen QA, and in Figure 13 for Trace QA.
Question: How many more seats does the Plaza Cinema have than it had when it opened?
Answer: 400
(a) The task is to retrieve and compute the difference between the number of seats mentioned in different context of the screen layout.
Question: How many days are there between the two gigs?
Answer: 16
(b) The task is to retrieve 2 dates and compute the difference between them in days. Note that the dates are in different colors of font and background, and in different date format. Also note the web page display show two images are missing, which is a common case the QA task is able to handle. We intentionally include screenshots as such.
Figure 10: Examples of Single Screen QA.
Question: How much do all the gym accessories cost in total?
Answer: \$127.77
Figure 11: An example of Multipage QA, about prices of selected gym accessories, each from a different online shop. Notice the layouts and font sizes vary among them.
Question: The sum of the prices of soup do jour, caviar hot dogs, and oysters do jour is how much more than the price of Tiramisu Della Casa?
Answer: \$59.5
Figure 12: An example of Multipage QA, about prices of food items over 4 websites of restaurant menus. Notice the layouts and font sizes vary significantly among them.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Question: To rent for an entire place, how much I can save per month if I rent a bedroom instead of two, if I rent the cheapest one?
Answer: \$400
Figure 13: An example of Trace QA. The example, situated in a rental exploration scenario, is about price differences between different options. Notice the trace captures the steps taken in the exploration.
Appendix B Statistics of the Dataset
We show detailed statistics of the dataset in Table 4, Table 4, and Table 4.
Category Num Examples
Restaurant 44
Shopping 184
Hotel 68
Transportation 114
Landmark 21
News 22
Media Sharing 7
Education & Information 9
Movies 24
Real Estate 19
Other 30
Total 542
Table 2: Distribution of the 542 examples of Single Screen QA, per categories and the number of screens in the example.
3 screens 4 screens 5 screens
54 90 27
Table 3: Distribution of the 171 examples of Multi Screen QA, per categories and the number of screens in the example.
Category Num Examples
Shopping 46
Restaurant 27
Media Sharing 8
News 55
Movies 14
Transportation 55
Real Estate 20
Hotel 12
Education & Information 20
Landmarks 10
Other 24
Total 292
Table 4: Distribution of examples of Trace QA for each category
Appendix C Prompts
In this section, we show the prompts used for evaluating the different MLLMs and we also discuss the improvement from prompt engineering.
C.1 Zero-shot prompt
You are given a sequence of screens and a question. Answer the question using only the information on the screens.
You should directly tell me your answer in the fewest words possible, and do not output any explanation or any other contents. Question:
C.2 Chain-of-thought prompt
You are given <num_screen> screenshots and a question. Your goal is to answer the question according to the screen information only.
Please follow the below steps to answer the question.
Question:
<question>
(Screenshots Analysis)
First, analyze the contents of each screenshot and list them here.
(Question Analysis)
What kind of information is needed from each screenshot to answer the question?
(Information Extraction)
Now, according to the question and analysis, please extract the relevant information from each screen and list them here.
(Analyze Information)
Based on the question, please analyze how to answer the question using the information above.
Please generate the answer with the information above. Please exclude any other additional words and information.
Appendix D Data Collection Details
D.1 Data Collection Instructions
The data collection process was performed with human raters who were responsible for capturing the screenshots and generating the corresponding questions and answers. In each task, raters are asked to collect screenshots and QA on a specific website category like Shopping, Restaurants etc. The instructions given to the raters for Multi Screen QA are shown below:
* Consider use cases where a user is researching a product to buy or comparing options from different web-sites.
* Capture 3-5 web screenshots taken from Google Chrome (desktop) spanning websites related to each other.
* Most of these screenshots should be from different websites.
* Generate a question-answer pair focusing on arithmetic or logical reasoning.
* Ensure that answering the question needs to use information from all the screenshots.
For the Trace QA split, similar to the Multi Screen QA split, the raters are asked to focus on specific website category. They are asked to install custom Chrome Extension which enables them to record their interactions on the screen. The chrome extension is meant to be turned on when the rater is ready to collect data. They can pause, review or delete any of the data captured. After collecting the trace, the raters add questions and answers related to the screenshots in the trace. The specific instructions are shown below:
* Consider use cases where a user is researching a product to buy or comparing options from different web-sites.
* Go to the first web page and start the trace recording.
* Navigate through the different webpages as part of the trace. Once the navigation is complete, click "stop" on the recording and review the screenshots captured.
* Delete any screenshots containing any personal information.
* Generate a question-answer pair focusing on arithmetic or logical reasoning.
* Ensure that answering the question needs to use information from about 3-5 screenshots.
D.2 Rater Details
For the Multi Screen QA task, we used 10 raters with ages between 25 and 35 based in India. The raters were employed as contractors to help with various data collection tasks apart from this task.
For the Trace QA task, the authors of the paper collected the data. They are based in the United States and Europe.